OpenStudy (anonymous):

Judah and 4 of his friends shared some apples. Each of the five friends had one and one-eighth of an apple. How many apples did they eat? four and four-ninths five and five-eighths five and one-eighth six and one-eighth

OpenStudy (anonymous):

9/8 x 5 = 45/8 = 5 5/8 The second one.